# Generative LiDAR Editing with Controllable Novel Object Layouts

### v1

@@ V-B Object insertion
The second part of our framework is object insertion. It is designed to allow users to specify the number, type, and pose of the objects to be inserted. In Fig. 7, we show some qualitative examples of object insertion.
However, the choice of object placement will also affect the quality of the generated data and matters if we want to evaluate the generated data quantitatively. Therefore, we design different object insertion strategies when we evaluate different aspects of the object insertion algorithm. Specifically, we have two experiments. The first reveals the domain gap between the generated data and the real data, and the second showcases the value of the generated data in improving downstream tasks like object detection.
@@ V-B1 Domain gap analysis
To rule out the impact of the choice of object placement on the realism of the generated data, we insert objects at the same pose as in the true Lidar scans. Specifically, we first remove all objects from the original dataset to obtain object-free Lidar scans. Then, we sample objects from the object library and insert them into the Lidar scan at the pose of the removed objects. In this way, we obtain a synthesized point cloud with the same pose layout of objects but with different objects. Note that while we do not change the pose layout, the pose of an inserted object is typically significantly different from the object’s original pose, making it a challenging object insertion task.
We evaluate the quality of the generated point clouds by running an object detection network VoxelNext [29] pretrained on nuScenes. The results are in Tab. II. As expected, the network performs differently on the true nuScenes dataset compared to our generated dataset. However, the results are comparable, with categories both outperforming and underperforming, suggesting a small domain gap between the two datasets.
V-B2 Value for downstream tasks
We also aim to demonstrate the practical value of our generated data in improving perception algorithms. To do this, we combine the real-world dataset with our generated dataset to fine-tune a detection network and evaluate its performance. In this experiment, we intentionally design our generated data to have different object layouts than the real data to introduce more variability into the training set. Specifically, we randomly perturb the target object poses by up to 2.5 meters and 45 degrees in yaw angle, while ensuring they remain in the ground area on the BEV map. The inserted objects are sampled from the object library. Tab. III shows that our generated data significantly improves the performance of the detection model. It shows that our synthetic data provides significant value in building more powerful algorithms.
@@ VI Conclusions
In this paper, we contribute a novel framework for LiDAR scene editing, generating realistic synthetic data with novel object layouts in a real-world environment. We leverage generative models to fill in the missing information about both the background and the objects, and adopt a spherical voxelization model to handle Lidar projection geometry efficiently. We demonstrate the effectiveness of our framework through experiments on a large-scale self-driving dataset, nuScenes, showing that the generated LiDAR point clouds closely resemble real-world data. We show incorporating these synthetic LiDAR scans as additional training data leads to significant improvements in object detection performance in real-world data. For future work, we plan to extend the object editing task beyond vehicle-type objects. We will also use novel view synthesis to build a flexible framework, where the novel-view information may come from temporarily continuous observations when available and from prior knowledge embedded in generative models otherwise.

### v2

@@ V-B Object insertion
The second part of our framework is object insertion. It is designed to allow users to freely specify the number, type, and pose of the objects to be inserted. In this experiment, we only manipulate the car category because AnchorFormer’s [27] point cloud completion is less reliable on other categories, but our framework is agnostic of the choice of completion models. In Fig. 7, we show some qualitative examples of object insertion.
The choice of object placement will also affect the quality of the generated data and matters if we want to evaluate the generated data quantitatively. Therefore, we design different object insertion strategies when we evaluate different aspects of the object insertion algorithm. Specifically, we have two experiments. The first reveals the domain gap between the generated synthetic data and the real data, and the second showcases the value of our synthetic data in improving autonomous driving downstream tasks such as object detection.
@@ V-B1 Domain gap analysis
To rule out the impact of the choice of object placement on the realism of the generated data, we insert objects at the same pose as in the true LiDAR scans. Specifically, we first remove all objects and inpaint the background. Then, we sample random objects from the object library and insert them into the LiDAR scan at the pose of the removed objects. In this way, we obtain a synthesized point cloud that maintains the original pose layout of objects while replacing the objects with different ones. We apply this procedure to the entire nuScenes validation set, producing a synthetic validation set of the same size as the original.
We assess the quality of our generated synthetic point clouds using the object detector VoxelNext [32], pretrained on nuScenes. We evaluate VoxelNext on both the (1) real and (2) synthetic nuScenes validation sets. In the first two rows of Tab. III, we can see that the mAP (mean average precision over all categories) remains similar when tested on the original validation set compared to our synthetic data, indicating a small domain gap. The car category’s higher AP shows that our inserted object point clouds have a small domain gap with the original ones, but they may also be too simple due to the lack of ray-drop modeling, indicating a direction for future improvements. The slightly lower APs of other categories are mostly due to the increased false positives caused by the background inpainting, which is helpful as we will see in Sec. V-B2.
V-B2 Values on autonomous driving downstream tasks
We aim to demonstrate the practical value of our generated data in improving perception algorithms. We compare two VoxelNext models, trained from scratch, using two different training procedures. The first model is trained only on real nuScenes data for 40 epochs. The second model is first pretrained on our synthetic training dataset (details provided below) for 15 epochs; then trained on real nuScenes data for an additional 40 epochs, following the same training schedule as the first model. In this experiment, we design our synthetic data to have different object layouts than the real data to introduce more variability. Specifically, we randomly perturb the target object poses by up to 2.5 meters and 45 degrees in yaw angle, while ensuring they remain in the ground area on the BEV map. The inserted objects are sampled from the object library. We apply this procedure to the entire nuScenes training set, generating a synthetic training set of the same size for pretraining. Tab. III shows that pretraining on our generated data (third row) yields superior performance compared to training exclusively on real nuScenes (second row). This shows that our synthetic data provides significant value in building more powerful algorithms. Interestingly, the improvement mostly comes from non-car categories by allowing the model to better suppress false positives. The inserted car objects, which are slightly easier to detect than the real ones, did not change the performance of the final model, but we expect it to improve with more detailed LiDAR modeling like ray-drop.
@@ VI Conclusions
This paper introduces a novel paradigm for LiDAR data generation by editing the object layouts in real-world LiDAR scans. Following this paradigm, our LiDAR-EDIT framework generates data with a small domain gap to the real world and gives users high controllability. Experiments show practical value of the generated data in downstream autonomous driving tasks. Incorporating ray-drop effects, optimizing object shape models, and automating the pose layout generation are interesting improving directions for future work.